# On the Federated Learning Framework for Cooperative Perception

### v3

@@ Abstract
Cooperative perception (CP) is essential to enhance the efficiency and safety of future transportation systems, requiring extensive data sharing among vehicles on the road, which raises significant privacy concerns. Federated learning offers a promising solution by enabling data privacy-preserving collaborative enhancements in perception, decision-making, and planning among connected and autonomous vehicles (CAVs). However, federated learning is impeded by significant challenges arising from data heterogeneity across diverse clients, potentially diminishing model accuracy and prolonging convergence periods. This study introduces a specialized federated learning framework for CP, termed the federated dynamic weighted aggregation (FedDWA) algorithm, facilitated by dynamic adjusting loss (DALoss) function. This framework employs dynamic client weighting to direct model convergence and integrates a novel loss function that utilizes Kullback-Leibler divergence (KLD) to counteract the detrimental effects of non-independently and identically distributed (Non-IID) and unbalanced data. Utilizing the BEV transformer as the primary model, our rigorous testing on FedBEVT dataset which is expanded on OpenV2V dataset, demonstrates significant improvements in the average intersection over union (IoU). These results highlight the substantial potential of our federated learning framework to address data heterogeneity challenges in CP, thereby enhancing the accuracy of environmental perception models and facilitating more robust and efficient collaborative learning solutions in the transportation sector.
@@ I Introduction
In a CP situation, dealing with data heterogeneity and distribution shifts is crucial. These challenges can significantly impact the performance and reliability of the system. For example, some clients may have the same type of sensors on one hand. However, their configurations such as mounting positions, angles, or calibration, may vary. On the other hand, some clients may have different types of sensors, which may lead to different data characteristics. Specifically, the dataset supplied by FedBEVT [27] incorporates data from three distinct client types: cars, buses, and trucks. Each type of vehicle has sensors installed at varying positions and angles, and the calibration among these sensors also differs, resulting in data heterogeneity. Moreover, environmental variations further influence data distribution shifts. For example, changes in lighting conditions across multiple time frames within a single scene can significantly alter. Such phenomena could be observed clearly in the dataset supplied by FedBEVT [27]. These are significant challenges for traditional federated learning methods, as it may be challenging to effectively synchronize the contributions of local models towards the convergence of the global model. This paper explores strategies to harness the benefits of external data while mitigating the impacts of data heterogeneity within a federated learning framework, with the aim of significantly improving the performance of local models.
@@
The rest of the paper are organised as follows: Section II provides related work, discussing CP, federated learning, and the application of federated learning within the CP context. Section III explains proposed FedDWA algorithm and DALoss function, with a formulation of the problem and describes the optimization objectives of our federated learning framework. In Section IV, our experimental setup and datasets are explained first, followed by details regarding applying FedDWA algorithm and DALoss function for BEV semantic segmentation task under CP setting. Finally, the conclusion is drawn in Section V.
@@ II-A Cooperative Perception in Autonomous Driving
It is essential that autonomous driving systems have the ability to perceive the environment accurately. This perception directly affects the efficiency of decision-making and overall safety. Traditional systems, constrained by limited sensing capabilities, often struggle with occlusions and distant objects. CP, enabled by V2X communications, marks a significant improvement, allowing CAVs to enhance perception through extensive wireless information sharing, encompassing V2V, V2I, and V2N communications. OpenCDA, introduced by [28], provides a versatile framework for developing and testing cooperative driving automation (CDA) systems within a simulated environment. Complementarily, the work [29] presented a large-scale simulated dataset for V2V perception, enriched with a diverse collection of scenes and annotated data. This dataset proves critical for advancing CP systems, as supported by [15], which underscores the expanded sensing capabilities facilitated by V2X. Further developments include the CoBEVT framework introduced in [30], a sparse vision transformer tailored for autonomous driving that leverages multi-agent multi-camera sensors, enabling collaborative perception across multiple vehicles. The work [31] discusses the integral role of advanced algorithms in optimizing CP systems. The works [32] and [33] explore the potential of integrating infrastructure-based sensors to bolster CP. There have been some notable advancements in perception frameworks recently. One of these is the introduction of V2X-ViT [34]. V2X-ViT is a novel vision Transformer designed to effectively combine information from on-road agents. Another important contribution is the work by Xu et al. [35]. They have created the first extensive real-world multimodal dataset for V2V perception. This dataset significantly enhances data availability for CP research. Lastly, Xiang et al. have introduced HM-ViT in their paper [36]. HM-ViT is an innovative multi-agent hetero-modal cooperative perception framework. It can accurately predict 3D objects in dynamic V2V scenarios. Collectively, these advancements represent a major leap towards more comprehensive and precise environmental perception in autonomous driving, addressing critical challenges and enhancing driving safety.
@@ II-C Federated Learning with Cooperative Perception
In this paper, we introduce an innovative federated learning framework tailored for V2X networks in CP, specifically addressing challenges posed by heterogeneity. Our framework is distinguished by its focus on the heterogeneity caused by Non-IID data across different traffic agents. We propose a novel loss function and an advanced aggregation algorithm designed to effectively manage the individual heterogeneity inherent in different agents, thereby enhancing the overall efficacy of model training and addressing the unique challenges of Non-IID data distribution.

### v4

@@ Abstract
Cooperative perception (CP) is essential to enhance the efficiency and safety of future transportation systems, requiring extensive data sharing among vehicles on the road, which raises significant privacy concerns. Federated learning offers a promising solution by enabling data privacy-preserving collaborative enhancements in perception, decision-making, and planning among connected and autonomous vehicles (CAVs). However, federated learning is impeded by significant challenges arising from data heterogeneity across diverse clients, potentially diminishing model accuracy and prolonging convergence periods. This study introduces a specialized federated learning framework for CP, termed the federated dynamic weighted aggregation (FedDWA) algorithm, facilitated by dynamic adjusting loss (DALoss) function. This framework employs dynamic client weighting to direct model convergence and integrates a novel loss function that utilizes Kullback-Leibler divergence (KLD) to counteract the detrimental effects of non-independently and identically distributed (Non-IID) and unbalanced data. Utilizing the BEV transformer as the primary model, our rigorous testing on FedBEVT dataset which is expanded on OpenV2V dataset, demonstrates significant improvements in the average intersection over union (IoU). These results highlight the substantial potential of our federated learning framework to address data heterogeneity challenges in CP, thereby enhancing the accuracy of perception models and facilitating more robust and efficient collaborative learning solutions in the transportation sector.
@@ I Introduction
In a CP situation, dealing with data heterogeneity and distribution shifts is crucial. These challenges can significantly impact the performance and reliability of the system. For example, some clients may have the same type of sensors on one hand. However, their configurations such as mounting positions, angles, or calibration, may vary. On the other hand, some clients may have different types of sensors, which may lead to different data characteristics. Specifically, the dataset supplied by FedBEVT [27] incorporates data from three distinct client types: cars, buses, and trucks. Each type of vehicle has sensors installed at varying positions and angles, and the calibration among these sensors also differs, resulting in data heterogeneity. Moreover, environmental variations further influence data distribution shifts. For example, changes in lighting conditions across multiple time frames within a single scene can significantly alter. Such phenomena could be observed clearly in the dataset supplied by FedBEVT [27]. These are significant challenges for traditional federated learning methods, as it may be challenging to effectively synchronize the contributions of local models towards the convergence of the global model. This paper explores strategies to harness the benefits of external data while mitigating the impacts of data heterogeneity within a federated learning framework, with the aim of significantly improving local models’ performance.
@@ II-A Cooperative Perception in Autonomous Driving
It is essential that autonomous driving systems have the ability to perceive the environment accurately. This perception directly affects the efficiency of decision-making and overall safety. Traditional single vehicle perception systems, although have recently been equipped by the BEV perception paradigm which extracts BEV space feature from input sequences in perspective view or 3D space, facilitate performance of perception tasks such as 3D bounding box detection and BEV segmentation [29, 28, 30], yet still often struggle with occlusions and distant objects since being constrained by limited sensing capabilities. CP, enabled by V2X communications, marks a significant improvement, allowing CAVs to enhance perception through extensive wireless information sharing, encompassing V2V, V2I, and V2N communications. OpenCDA, introduced by [32], provides a versatile framework for developing and testing cooperative driving automation (CDA) systems within a simulated environment. Complementarily, the work [33] presented a large-scale simulated dataset for V2V perception, enriched with a diverse collection of scenes and annotated data. This dataset proves critical for advancing CP systems, as supported by [15], which underscores the expanded sensing capabilities facilitated by V2X. Further developments include the CoBEVT framework introduced in [31], a sparse vision transformer tailored for autonomous driving that leverages multi-agent multi-camera sensors, enabling collaborative perception across multiple vehicles. The work [34] discusses the integral role of advanced algorithms in optimizing CP systems. The works [35] and [36] explore the potential of integrating infrastructure-based sensors to bolster CP. There have been some notable advancements in perception frameworks recently. One of these is the introduction of V2X-ViT [37]. V2X-ViT is a novel vision Transformer designed to effectively combine information from on-road agents. Another important contribution is the work by Xu et al. [38]. They have created the first extensive real-world multimodal dataset for V2V perception. This dataset significantly enhances data availability for CP research. Lastly, Xiang et al. have introduced HM-ViT in their paper [39]. HM-ViT is an innovative multi-agent hetero-modal cooperative perception framework. It can accurately predict 3D objects in dynamic V2V scenarios. Collectively, these advancements represent a major leap towards more comprehensive and precise environmental perception in autonomous driving, addressing critical challenges and enhancing driving safety.